OpenStudy (anonymous):

Given f(x) = −15x − 2x2 + x3, is there a zero of f(x) between a = 4 and b = 6?

OpenStudy (anonymous):

looks like there is because f(4) is negative and f(6) is positive

OpenStudy (anonymous):

lets try factoring this - x is the gcf: = x(x^2 - 2x - 15) = x (x - 5)(x + 3) right - that confirms your query - one zero is between 4 and 6 (that is 5 )and the other two are 0 and -3